OpenStudy (anonymous):

Solve for x: 3x-1/5 - 2x-4/3 = 1

OpenStudy (luigi0210):

Get common denominators

OpenStudy (anonymous):

15?

OpenStudy (luigi0210):

.. Or cancel them out

OpenStudy (luigi0210):

Yes, now multiply everything by 15 I thought you were better than that Nurali

OpenStudy (anonymous):

That wasn't it, @Nurali, its a whole number. But thanks.:)

OpenStudy (anonymous):

Everything? Including numerators and denominators?

OpenStudy (anonymous):

Do I plug 15 into x?

OpenStudy (luigi0210):

\[15[\frac{ 3x-1 }{ 5 }-\frac{ 2x-4 }{ 3 }]=1*15\]

OpenStudy (anonymous):

I multiplied everything but I'm still a little confused.

OpenStudy (nurali):

\[\frac{ 3x-1 }{ 5 }-\frac{ 2x-4 }{ 3 }=1\] \[\frac{ 3(3x-1)-5(2x-4) }{ 5*3 }=1\] \[\frac{ 9x-3-10x+20 }{ 15 }=1\] \[\frac{ -x+17 }{ 15 }=1\] \[-x+17=15\] \[x=17-15\] \[x=2\]

OpenStudy (luigi0210):

.. Thank you Nurali

OpenStudy (anonymous):

Okay good that's what I got. Thank you both!

OpenStudy (nurali):

My Pleasure.